Hii... Any one of you guys can explain about dot product and cross product of vectors. And please suggest good books for basic concepts of physics at IIT level.

Grade:11

3 Answers

Nilesh
8 years ago
sa
erra akhil
8 years ago
Dear     ,
DOT PRODUCT of 2 vectors.
1.)It gives a sclar.
2.)Magnitude of dot product is equal to product of magnitude of two vectors and the cosine of angle between them.
Example:- Let A and B be two vectors with angle between them equal to θ.
A.B=(magnitude of A)(Magnitude of B)(cosθ)
erra akhil
8 years ago
Cross Product:-
1.)Vector product of 2 vectors gives a vector.
2.)The resultant product vector of 2 vectors is in the direction perpendicular to the plane in which the given vectors are present.
3.)Magnitude of cross product of  2 vector is the product of magnitude of 2 vectors and the sine of angle between them.
Example:-
If A and  B are two vectors, 
A = ai + a2 j + a3 k
B = b1 i + b2 j + b3 k
magnitude of A*B=(magnitude of A)(Magnitude of B)(sin θ)
A*B=
| i        j      k |
| a1     a2    a|
| b1    b2    b3|
i,j,k are components of A and B along x,y,z axes
Approve if my answer helped you